OpenStudy (anonymous):

The width of a rectangle is 12 units. Can the perimeter x of the rectangle can be 60 units when its length y is 18 units?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok i need help choosing the questions? The width of a rectangle is 12 units. Can the perimeter x of the rectangle can be 60 units when its length y is 18 units? (4 points) No, the rectangle cannot have x = 60 and y = 18 because x + y is less than 60. No, the rectangle cannot have x = 60 and y = 18 because x + y is less than 24. Yes, the rectangle can have x = 60 and y = 18 because x = 12 + 2y. Yes, the rectangle can have x = 60 and y = 18 because x = 24 + 2y. @WinterRaine

OpenStudy (anonymous):

Looks like the answer is: Yes, the rectangle can have x = 60 and y = 18 because x = 24 + 2y.

OpenStudy (anonymous):

ok thank you! do you mind helping me with two more problems? @WinterRaine

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

Are you going to post them separate or in this question?

OpenStudy (anonymous):

Im going to post them here. @WinterRaine

OpenStudy (anonymous):

The table below shows two equations: Equation 1 |3x - 1| + 7 = 2 Equation 2 |2x + 1| + 4 = 3 Which statement is true about the solution to the two equations? (4 points) Equation 1 and equation 2 have no solutions. Equation 1 has no solution and equation 2 has solutions x = 0, 1. The solutions to equation 1 are x = -1.3, 2 and equation 2 has no solution. The solutions to equation 1 are x = -1.3, 2 and equation 2 has solutions x = 0, 1.

OpenStudy (anonymous):

I'd say that they both have no solutions because since it's multiple choice, all you have to do is plug in your x's. Make sense?

OpenStudy (anonymous):

yeah but which one do i choose? which answer?

OpenStudy (anonymous):

The no solutions. First one.

OpenStudy (anonymous):

Ok this is the second one I need help with @WinterRaine Below are the steps to solve an equation: Step 1: |x - 2| + 3 = 7 Step 2: |x - 2| = 7 - 3 Step 3: |x - 2| = 4 Which of the following is a correct next step to solve the equation? (4 points) x + 2 = -4 -x - 2 = 4 x + 2 = 4 x - 2 = -4

OpenStudy (anonymous):

the second one is the closest

OpenStudy (anonymous):

it had to be either -x - 2=4 OR x - 2=4